Two pieces of clay are moving directly toward each other. When they collide, they stick together and move as one piece. One piece having mass 300~\text{g}300 g is moving to the right at a speed of 1~\text{m/s}1 m/s. The other piece has mass 600~\text{g}600 g and is moving to the left at a speed of 0.75~\text{m/s}0.75 m/s. Use momentum conservation to determine the velocity of the combined clay piece after they collide. (Note: you can treat this as a one-dimensional problem, but since you're computing velocity not speed, you need to include an appropriate sign depending on whether the final piece is moving to the right or left)
Answer #1

Let the speed of combined clay particle be V m/s

by the conservation of momentum we have

m1u1 + m2u2 = (m1 + m2) V

300*1 + 600*-0.75 = (300 + 600)*V

V = -0.167 m/s

so the combined clay particle moves to the left with 0.167 m/s